OpenStudy (anonymous):

2x^2 = 8x y^2/7 = y Solve for x on the first, and solve for y on the other. They're separate problems. I need to study for a test, and don't know how to do these, so if you could show work, that'd be great!

OpenStudy (anonymous):

So. we have this: \[2x^{2}=8x\] Now substract 8x from both sides and we will have \[2x^{2}-8x=8x-8x\]: \[2^{2}x-8x=0\]

OpenStudy (anonymous):

So far I'm understanding it. so what's x? Or do we substitute 0 for x

OpenStudy (anonymous):

This is has the form: \[ax^{2}+bx+c=0\] where a is the coefficient that multiplies x^2, b is the coefficient of x and c is the constant

OpenStudy (anonymous):

And you can solve it with the quadratic formula: \[x=\frac{ -b \pm \sqrt{b^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

It shouldn't be that complicated.

OpenStudy (anonymous):

No it is not

OpenStudy (anonymous):

\[2x^2-8x=0,2x(x-4)=0\] either x=0, or x-4=0 x=4

OpenStudy (anonymous):

Oh so you basically "foiled it" I see.

OpenStudy (anonymous):

\[\frac{ y^2 }{7 }=y\]

OpenStudy (anonymous):

Factored

OpenStudy (anonymous):

multiply by 7 and solve like above question.